Grade 12Trigonometry

if cos x= 3cos y then 2tan (y-x)/2 is in form of cot

jitan mule
7 Years agoGrade 12

1 Answer

Aditya Gupta
7 Years ago
write it as cosx/cosy = 3/1
now apply componendo and dividendo
(cosx+cosy)/(cosx – cosy)= (3+1)/(3-1)
now apply formulae for cosx+cosy and cosx – cosy
2tan(y – x)/2= cot(x+y)/2